Question 4: The greatest movie of all time, John Carter (2012), was one of the biggest flops in movie making. The movie lost $200 million dollars and that was after all the revenue was earned! I guess a sequel might be out of the question. I can say one thing and that is Clear Eyes, Full Hearts, Sometimes Lose J

Anyways back to the problem. Let us pretend a sample of 40 theaters that showed the movie averaged ticket sales revenue of $25,000 and a standard deviation of $4000.

  1. Find the 95% confidence interval for the true average sales revenue per theater
  2. Using the movie ticket price of $9.50, what is the 95% confidence interval for the mean number of customers per theater
Answer #1

Ans:

a)As, n>=30,we can use normal distribution.

95% confidence interval for the true average sales revenue per theater

=25000+/-1.96*(4000/sqrt(40))

=25000+/-1240

=(23760, 26240)

b)

sample mean for number of customers per theater=25000/9.50=2632

standard deviation=4000/9.5=421

95% confidence interval for mean number of customers per theater

=2632+/-1.96*(421/sqrt(40))

=2632+/-130

=(2502, 2762)